### v1

# Symphony-MoE: Harmonizing Disparate Pre-trained Models into a Coherent Mixture-of-Experts

Abstract
Mixture-of-Experts (MoE) models enable scalable performance by activating large parameter sets sparsely, minimizing computational overhead. To circumvent the prohibitive cost of training MoEs from scratch, recent work employs upcycling, reusing a single pre-trained dense model by replicating its feed-forward network (FFN) layers into experts. However, this limits expert diversity, as all experts originate from a single pre-trained dense model. This paper addresses this limitation by constructing powerful MoE models using experts sourced from multiple identically-architected but disparate pre-trained models (e.g., Llama2-Chat and Code Llama). A key challenge lies in the fact that these source models occupy disparate, dissonant regions of the parameter space, making direct upcycling prone to severe performance degradation. To overcome this, we propose Symphony-MoE, a novel two-stage framework designed to harmonize these models into a single, coherent expert mixture. First, we establish this harmony in a training-free manner: we construct a shared backbone via a layer-aware fusion strategy and, crucially, alleviate parameter misalignment among experts using activation-based functional alignment. Subsequently, a single lightweight stage of router training coordinates the entire architecture. Experiments demonstrate that our method successfully integrates experts from heterogeneous sources, achieving an MoE model that significantly surpasses baselines in multi-domain tasks and out-of-distribution generalization.
1 Introduction
Figure 1: Comparison between the workflow of naive upcycling and ours.
Large language models (LLMs) have demonstrated remarkable progress in recent years, with the Mixture-of-Experts (MoE) architecture emerging as a key paradigm for efficient scaling (Shazeer et al. 2017). By dynamically routing computations to sparse subsets of parameters—known as “experts”—MoE models achieve vast total capacity while maintaining a near-constant floating-point operations (FLOPs) count per forward pass.
Despite their potential, training MoE models from scratch requires prohibitive computational resources and massive datasets. To mitigate this, academia and industry have explored upcycling techniques (Komatsuzaki et al. 2022; Zhu et al. 2024). These methods typically expand a pre-trained dense model’s feed-forward network (FFN) layers into MoE layers, initializing each expert as a replica of the original FFN. Subsequent continual training then encourages expert specialization, with studies confirming that diverse data subsets can effectively differentiate replicated experts (Sukhbaatar et al. 2024).
However, existing upcycling approaches such as “Branch-Train-Mix” (Sukhbaatar et al. 2024) share a fundamental limitation: reliance on a single pre-trained checkpoint (see Figure 1). Consequently, all experts originate from identical parameter space, regardless of subsequent specialization. This “single-origin” constraint intrinsically caps expert diversity. In practice, a large number of high-quality, identically architected models—specialized through distinct pre-training objectives, data corpora, or versions—exist, such as the dialogue-optimized Llama2-Chat (Touvron et al. 2023) and the code-centric Code Llama (Roziere et al. 2023). These models capture unique knowledge and capabilities within their domains, occupying valuable yet heterogeneous regions of the parameter space. Directly leveraging such specialized models as distinct experts offers a promising avenue for constructing more powerful and versatile MoEs, but remains infeasible under current frameworks.
Figure 2: Overview of the Symphony-MoE construction pipeline. In Stage 1, disparate source models are harmonized without training: non-FFN layers are merged into a shared backbone using techniques such as SLERP, while FFN layers are aligned through activation-based neuron permutation. In Stage 2, only the router is trained while keeping the backbone and expert parameters frozen, enabling coordination among the now-compatible experts.
Thus, this paper addresses a core research question: Can we effectively assemble MoE models by upcycling experts from disparately initialized source models? Naive upcycling encounters severe parameter space misalignment (Ainsworth, Hayase, and Srinivasa 2022). Models trained with divergent histories tend to occupy incompatible regions of both the numerical and semantic parameter space. Simply concatenating such models leads to catastrophic interference, as the misaligned experts disrupt each other’s internal representations. As a result, the router can no longer perform meaningful dispatching across functionally incoherent experts.
We propose Symphony-MoE, a framework for harmonizing a collection of disparate models into a single, powerful MoE. We achieve this harmonization through a two-stage process, where the parameter space misalignment is resolved via training-free functional alignment. In this first stage, we construct a coherent shared backbone using layer-aware fusion. We apply Spherical Linear Interpolation (SLERP) (Wortsman et al. 2022) to self-attention modules to maintain their geometric integrity, while using a selective linear strategy for token embeddings to handle vocabulary mismatches and simple averaging for the statistically-driven LayerNorm parameters. Crucially, to harmonize the experts themselves, we functionally align them to a shared representational space. This is achieved not by averaging, but by a precise reordering of neurons based on activation similarities, solved via a Linear Assignment Problem. This training-free process yields a pool of diverse yet functionally compatible experts, setting the stage for efficient coordination.
The second stage implements lightweight global coordination. We freeze the expert layers and shared backbone to train the router independently. To prevent expert collapse and ensure balanced utilization, this stage critically employs an auxiliary loss function promoting load balancing across experts.
The core contributions of this work are threefold:
(1) We investigate the problem of upcycling multiple identically architected yet independently pretrained models into a unified Mixture-of-Experts (MoE) system. We identify parameter space misalignment as a key barrier to this process..
(2) We introduce a lightweight framework that addresses expert incompatibility without gradient-based training, combining layer-aware fusion and activation-driven functional alignment to preserve the specialized knowledge embedded in each source model.
(3) Through comprehensive experiments on in-distribution and out-of-distribution datasets, we show that our method achieves strong performance and generalizes well to out-of-distribution settings, demonstrating the feasibility of upcycling disparate pre-trained models into unified expert systems.
2 Related Works
2.1 Mixture-of-Experts Models
The Mixture-of-Experts (MoE) paradigm is originally proposed as a “divide and conquer” modular learning system (Jacobs et al. 1991). Its modern renaissance was triggered by the Sparsely-Gated MoE (Shazeer et al. 2017), which introduced sparse, top-k routing to enable conditional computation at an unprecedented scale. This decouples the model’s parameter count from its computational cost.
Subsequent advances, most notably the Switch Transformer (Fedus, Zoph, and Shazeer 2022), simplified the routing mechanism to top-1 selection, improving training stability and reducing communication overhead. This streamlined approach established sparsity as a core strategy for scaling, as adopted in models like GLaM (Du et al. 2022) and Mixtral (Jiang et al. 2024). However, training these massive MoE models from scratch remains prohibitively expensive, which motivates more resource-efficient construction methods.
2.2 MoE Construction via Upcycling
To mitigate the immense cost of training MoEs from scratch, the upcycling paradigm has emerged as a practical alternative. The vanilla approach, or Sparse Upcycling (Komatsuzaki et al. 2022; He et al. 2024), involves converting a pre-trained dense model into an MoE by replicating its feed-forward network (FFN) layers to form multiple, initially identical experts. While this provides a strong initialization, it suffers from a knowledge bottleneck and a lack of initial expert diversity, as all experts originate from a single source.
To address this, Drop-Upcycling (Nakamura et al. 2025) performs partial re-initialization to enhance expert diversity. “Branch-Train-Mix” strategy (Sukhbaatar et al. 2024; Horoi et al. 2025; Kang et al. 2024; Li et al. 2024a) constructs an MoE from multiple, pre-specialized dense models to inject diverse knowledge from the outset. For example, BAM (Zhang et al. 2024) first fine-tunes copies of a base model on different domains and then uses their specialized FFN and attention layers as the experts in the final MoE. However, it introduces a new critical challenge: how to effectively fuse the knowledge from these disparate trained specialists, whose parameters reside in different regions of the loss landscape. Current methods like BAM and Self-MoE (Kang et al. 2024) rely on rudimentary fusion techniques, such as simple averaging for non-expert layers, which are not guaranteed to be optimal.
2.3 Model Merging
The challenge of combining disparate pre-trained models is the central focus of the model merging field (Tao et al. 2024). To merge truly “stranger” models, advanced techniques explicitly resolve misalignment. Model Soups (Wortsman et al. 2022) averages the weights of models fine-tuned from the same checkpoint, assuming they remain in the same well-connected loss basin. Methods like PLeaS (Ito, Yamada, and Kumagai 2024) and Git Re-Basin (Ainsworth, Hayase, and Srinivasa 2022) find an optimal permutation matrix to align neurons layer-wise before averaging. TIES-Merging (Yadav et al. 2023) resolves interference between task vectors by trimming redundant updates and electing a dominant sign for parameter changes. These methods, however, are general-purpose algorithms designed to produce a single, dense consensus model. They are not tailored for the specific architectural needs of MoE construction, such as preserving expert diversity or informing the initialization of the router.
Our work bridges this gap. We integrate permutation-based alignment directly into the MoE construction pipeline, repurposing these techniques for a new architectural goal. We use alignment to make disparate experts functionally compatible. This ensures they can operate within a shared coordinate system while preserving their individual specializations.
3 Methods
We introduce a two-stage framework to effectively assemble a MoE model from disparate, pre-trained source models. The first stage operates entirely without training, resolving the critical parameter space misalignment problem through training-free functional alignment. The second stage then trains a router that coordinates the now-compatible experts, unlocking their collective capabilities.
3.1 Preliminaries and Problem Formulation
We first establish the requisite notation for Transformer and MoE architectures and then formally define the problem.
A standard Transformer model is composed of stacked layers, each typically containing a multi-head self-attention mechanism followed by a position-wise feed-forward network (FFN). In an MoE architecture, the dense FFN layer is replaced by a set of $N$ independent “expert” networks $\{E_{1},E_{2},\ldots,E_{N}\}$, each with the same architecture as the original FFN. A trainable router network, $R$, directs the input tokens to a sparse subset of these experts. For each input token $x$, the router computes gating values $g(x)$ that determine how the outputs of the experts are combined:
$M_{MoE}(x)=\sum_{i=1}^{N}g_{i}(x)\cdot E_{i}(x)$ (1)
The gating values are typically produced by a linear layer followed by a softmax function, $g(x)=Softmax(xW_{g})$, where $W_{g}$ is a trainable routing matrix. For sparsity, often only the top-k experts with the highest gating values are activated for any given token.
Our work addresses the novel challenge of constructing an MoE model from multiple, disparate pre-trained source models. Formally, we are given a set of $N$ source models, $\{M_{1},M_{2},\ldots,M_{N}\}$. These models share an identical architecture but possess distinct sets of parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ as a result of being trained on different datasets $\{D_{1},D_{2},\ldots,D_{N}\}$ , for different tasks, or with different optimization objectives. The goal is to construct a single, powerful MoE model, $M_{MoE}$, by leveraging the FFN layers of the source models as its experts. The fundamental barrier to this goal is the severe parameter space misalignment between the models, where the parameters $\theta_{i}$ and $\theta_{j}$ (for $i\neq j$) occupy incompatible numerical and semantic spaces.
3.2 Stage 1: Training-Free Functional Alignment
Our first stage assembles a cohesive MoE model from diverse parent models entirely without training. This process involves two key steps: fusing the shared architectural backbone and functionally aligning the FFN layers, which will serve as the experts.
Shared Backbone Construction. We first construct the shared, non-expert layers. Our approach to this fusion is explicitly layer-aware. Rather than applying a single, uniform merging technique, we tailor the fusion strategy to the specific architectural function of each component. Specifically:
$\cdot$ For the token embedding layers, we follow the selective linear strategy implemented in MergeKit (Goddard et al. 2024). For tokens shared across models, the algorithm retrieves their corresponding embedding vectors from each model and performs standard linear averaging. For tokens that are unique to a specific model—i.e., not present in the vocabularies of others—the original embedding is preserved without modification.
$\cdot$ For self-attention modules, which are critical for contextual understanding, we apply Spherical Linear Interpolation (SLERP) (Wortsman et al. 2022). This method better preserves the geometric integrity of the weight space and mitigates the functional degradation often caused by naive linear averaging. We apply it to all weight matrices (Q, K, V, O projections).
$\cdot$ In contrast, the primary statistical function of LayerNorm parameters makes a simple linear average a sufficient and stable choice  (Jin et al. 2022).
Pairwise Inter-Expert Alignment. We designate one model’s FFNs as the “anchor” and then permute the neurons of the other models’ FFNs to match the anchor’s functional behavior based on activation patterns.
For each non-anchor model $M_{i}$ and each layer $l$, the alignment process is as follows:
1. Activation Collection: We construct a small and diverse calibration set $D_{\mathrm{cal}}$ by sampling 5% from each instruction fine-tuning dataset $D_{1},D_{2},\ldots,D_{N}$} used in dense model training. Each instruction in $D_{\mathrm{cal}}$ is passed through both the anchor model $M_{1}$ and a target model $M_{i}$ to extract their FFN activation matrices at layer $l$, denoted as $A_{1}^{(l)}$ and $A_{i}^{(l)}$, respectively.
2. Permutation Matching: We find the optimal permutation matrix $P_{i}^{(l)}$ that aligns the neurons of $M_{i}$ to $M_{1}$ by solving the linear assignment problem:
[…]
where $\mathcal{P}$ is the set of permutation matrices and $||\cdot||_{F}$ is the Frobenius norm. This problem can be solved efficiently with the Hungarian algorithm (Mills-Tettey, Stentz, and Dias 2007).
3. Weight Remapping: We apply the computed permutation $P_{i}^{(l)}$ to the FFN weights (Nasery et al. 2025). This involves permuting the output dimension of the first linear layer’s weights $W_{up}$ and the input dimension of the second’s $W_{down}$ to maintain functional equivalence:
$W^{\prime}_{up,i}=W_{up,i}^{(l)}P_{i}^{(l)},\quad W^{\prime}_{down,i}=(P_{i}^{(l)})^{T}W_{down,i}^{(l)}$ (3)
This procedure results in $N$ functionally aligned experts for each layer, all operating in a consistent parameter space.
We analyzed the computational cost to verify the scalability. The analysis can be found in the Appendix.
3.3 Stage 2: Activating the MoE via Lightweight Router Training
In Stage 2, we integrate the constructed shared backbone with the expert models and introduce a randomly initialized router, implemented as a simple linear layer. The expert layers and shared backbone—which encode knowledge from the source models—are frozen, and only the router is trained. The calibration dataset $D_{\mathrm{cal}}$, introduced in Stage 1, is reused for this purpose, enabling the router to learn how to coordinate the experts.
We adopt a top-2 routing mechanism and train the router using a composite loss function:
$\mathcal{L}_{total}=\mathcal{L}_{lm}+\lambda\cdot\mathcal{L}_{bal}$ (4)
where $\mathcal{L}_{lm}$ is the standard causal language modeling objective, and $\mathcal{L}_{bal}$ is a load-balancing regularization term. The coefficient $\lambda$ is fixed at 0.01 in all experiments. $\mathcal{L}_{bal}$ encourages uniform expert utilization, which mitigates expert collapse and allows the model to fully leverage the diversity of its specialists.
Notably, the router is not provided with explicit supervision indicating which expert to select for each token, as there is no ground-truth mapping between tokens and experts. Instead, it learns expert assignment implicitly through training on instruction–response pairs in $D_{\mathrm{cal}}$.
4 Experiments
In this section, we aim to address the following key research questions: (1) How does Symphony-MoE perform at different model scales or types, compared to baseline models on both in-distribution and out-of-distribution (OOD) datasets? (2) Does our proposed alignment method effectively mitigate parameter space misalignment across experts? (3) How does the choice of anchor model influence the final performance of Symphony-MoE? (4) What is the impact of varying the number of experts on model performance? (5) Ablation Analysis: How sensitive is the system to variations in alignment and merging strategies?
4.1 Experimental Setup
4.1.1 Build Disparately Initialized Dense Models
ID OOD
MMLU GSM8K BBH HumanEval TruthfulQA Avg. MedCQA
Dense (0.5B)
General ($M_{1}$, ANC) 44.27±0.40 18.87±1.07 27.46±0.50 24.40±3.36 27.29±1.56 28.46 26.20±0.19
Math ($M_{2}$) 42.07±0.40 19.33±1.08 22.72±0.46 20.73±3.17 28.40±1.58 26.65 26.23±0.19
Code ($M_{3}$) 30.96±0.38 20.48±0.41 25.33±0.49 28.71±3.27 25.81±1.55 26.26 25.93±0.19
Science ($M_{4}$) 39.25±0.41 21.30±1.12 22.98±0.48 18.29±3.02 30.01±1.50 26.37 25.56±0.19
MoE (0.5B $\times$ 4)
BTX 37.28±0.61 18.95±0.42 25.57±0.48 26.59±3.15 23.90±1.19 26.46 24.02±0.19
BAM 42.76±0.70 19.15±0.68 27.68±0.58 26.77±3.07 26.54±1.61 28.58 26.79±0.19
Drop 44.08±0.40 22.85±0.39 26.51±0.48 25.52±3.19 30.81±1.55 29.95 27.95±0.20
Symphony (Ours) 45.10±0.40 24.57±1.11 29.64±0.48 28.02±3.12 31.54±1.50 31.77 29.07±0.19
Table 1: Performance comparison of dense (Qwen2/2.5 0.5B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data. The best and second-best results are bolded and underlined. “ANC” stands for anchor model.
ID OOD
MMLU GSM8K BBH HumanEval TruthfulQA Avg. MedCQA
Dense (1.5B)
General ($M_{1}$, ANC) 59.57±0.39 33.72±0.39 44.39±0.55 37.19±3.79 29.87±1.60 40.95 30.39±0.20
Math ($M_{2}$) 56.73±0.39 34.11±0.50 43.97±0.54 37.20±3.79 31.33±1.62 40.67 30.36±0.20
Code ($M_{3}$) 43.74±0.41 35.86±1.33 36.17±0.55 43.41±3.81 26.19±1.51 37.07 28.72±0.19
Science ($M_{4}$) 54.60±0.40 32.45±1.30 36.05±0.53 35.37±3.74 30.78±1.57 37.85 29.72±0.19
MoE (1.5B $\times$ 4)
BTX 45.12±0.88 30.35±0.45 40.11±0.58 29.08±3.15 25.02±1.15 33.94 26.92±0.18
BAM 50.77±0.75 36.99±0.61 45.01±0.59 37.84±2.89 27.14±1.89 39.55 28.97±0.18
Drop 57.14±0.81 34.92±1.12 44.83 ±0.43 36.72±1.89 31.88±1.49 41.10 32.90±0.19
Symphony (Ours) 58.91±0.31 39.12±1.09 46.97±0.36 42.39±3.69 32.95±2.31 44.07 35.26±0.22
Table 2: Performance comparison of dense (Qwen2/2.5 1.5B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data. The best and second-best results are bolded and underlined. “ANC” stands for anchor model.
To ensure a rigorous evaluation of our proposed merging method, we first constructed a set of four disparately initialized dense models at scale of 0.5/1.5B. This experimental suite was designed to reflect real-world disparity by systematically incorporating variations across three dimensions: model versions, pre-training data domains, and downstream instruction-tuning tasks.
Our framework is built upon three distinct foundation checkpoints, each contributing a different source of disparity. The first is Qwen2.5-Base, a general-purpose model trained on broad-domain natural language data, with a focus on tasks involving understanding and generation. To introduce task- and data-level disparity, we include Qwen2.5-Coder (Hui et al. 2024), which is primarily trained on code-related corpora and optimized for tasks such as code completion and code–natural language interconversion. Finally, to incorporate version-level disparity, we use Qwen2-Base (Team 2024), an architecturally compatible predecessor of Qwen2.5-Base with a completely different pre-training history.
Based on the three foundational models, we derived four specialist experts through instruction fine-tuning. The generalist expert ($M_{1}$) and the mathematics expert ($M_{2}$) were obtained by fine-tuning Qwen2.5-Base on the Alpaca (Maeng, Colin, and Lucia 2017) and MetaMathQA (Yu et al. 2024) datasets, respectively. The code expert ($M_{3}$) was derived from our code-specialized foundation using the CodeAlpaca (Li et al. 2024b) dataset. The science expert ($M_{4}$) was created by fine-tuning Qwen2-Base on the SciQAG (Wan et al. 2024) dataset.
Further training details are provided in the Appendix.
4.1.2 Evaluation Datasets and Metrics
We evaluate our method on five in-distribution datasets: MMLU (General)  (Hendrycks et al. 2021), GSM8K (Math) (Cobbe et al. 2021), BBH (General, Reasoning)  (Suzgun et al. 2022), HumanEval (Code) (Chen et al. 2021), and TruthfulQA (Science) (Lin, Hilton, and Evans 2022). We also include an out-of-distribution benchmark, MedCQA (Medicine) (Shoham and Rappoport 2024).
For evaluation, we follow the standard metrics used in each benchmark: MMLU and MedCQA are evaluated using zero-shot accuracy, GSM8K and BBH use eight-shot accuracy, HumanEval is measured with pass@1, and TruthfulQA adopts the MC1 metric.
4.1.3 Baselines
We compare our upcycling approach against two categories: the dense models $M_{1}$, $M_{2}$, $M_{3}$, $M_{4}$ that we upcycle, and recent upcycling methods (BTX (Sukhbaatar et al. 2024), BAM (Zhang et al. 2024), Drop (Nakamura et al. 2025)). To ensure fairness, the baseline method uses the same data as ours for fine-tuning the dense model. We reproduced the results according to the original paper, and the reproduction details are in the appendix.
4.2 Implementation Details
Figure 3: Quantitative analysis of inter-expert functional specialization using Centered Kernel Alignment (CKA). Lower CKA scores indicate greater functional specialization. High CKA scores between experts reflect parameter space misalignment, caused by failing to align neurons functionally during merging. This misalignment leads to representational collapse, erasing the distinct capabilities of individual experts.
Our experiments were conducted on 8 NVIDIA V100 GPUs. The dense model was implemented using LLaMA-Factory (Zheng et al. 2024), and trained for for 2 epoches. Detailed training configurations are provided in the Appendix.
In Stage 1: Alignment, we randomly sampled from each training dataset to construct the calibration set $D_{\mathrm{cal}}$, resulting in a total of 10.404 million tokens. To ensure consistent alignment and integration across expert models, we fixed a general-purpose expert model $M_{1}$ as the anchor model. To ensure that the knowledge of each expert is equally integrated and to avoid introducing any bias, the weights for merging the 4 dense models are set to 0.25.
In Stage 2: Router Training, we set the cutoff length to 2048 and train the model for 6 epochs with a batch size of 2. The learning rate is fixed at 5e-5, and optimization is performed using AdamW (Zhuang et al. 2022) with a maximum gradient norm of 1.0.
In the evaluation phase, we used lm-evaluation-harness-v0.4 (Gao et al. 2024) to obtain scores for datasets such as MMLU and GSM8K.
4.3 Main Results
Overall Performance. As summarized in Table 1 and 2, Symphony-MoE achieves the highest average score across all evaluation datasets, substantially outperforming all MoE baselines, including BTX, BAM, and Drop. In contrast, individual dense models demonstrate strong domain specialization but suffer severe performance degradation on out-of-domain tasks, indicating limited generalization.
In-Domain Expertise Preservation. Importantly, Symphony-MoE is not a naive average of its experts. On in-domain tasks, it effectively retains the specialized capabilities of its constituent models. For example, in the HumanEval, although the dedicated code expert $M_{3}$ achieves the top score, Symphony-MoE performs comparably, trailing by only 1–2 percentage points.
Out-of-Distribution Generalization. Beyond domain retention, Symphony-MoE demonstrates superior generalization ability, outperforming all baselines on MedCQA. This suggests that our activation-based alignment strategy enables the model to internalize more transferable and fundamental reasoning patterns, rather than merely interpolating between expert skills.
Robustness Across Scales. We further validate these findings across two model scales—4 experts × 0.5B and 4 experts × 1.5B. In both settings, Symphony-MoE consistently exhibits strong average performance, competitive in-domain expertise, and leading out-of-distribution generalization, confirming the robustness of our approach.
Robustness Across Types. We further validated the effectiveness of Symphony-MoE using the Llama model as the source architecture. As shown in the Appendix, our method yields consistent improvements, demonstrating its robustness across different backbone model types.
4.4 Analysis of Inter-Expert Functional Specialization
To quantitatively assess the effectiveness of our functional alignment method, we compute the pairwise functional specialization between expert layers using Centered Kernel Alignment (CKA) (Cortes, Mohri, and Rostamizadeh 2012), a widely adopted technique for comparing neural network representations. A lower CKA score indicates greater functional specialization. We evaluate three scenarios: (a) the original, unmerged expert models as a reference baseline; (b) a naive merging approach without alignment; and (c) our proposed method employing activation-based functional alignment.
As shown in Figure 3, the naive merging strategy leads to a severe representational collapse. Experts trained independently develop distinct internal coordinate systems; merging them without alignment forces the model to conflate functionally unrelated neurons. This results in inflated inter-expert CKA scores (e.g., 0.65–0.75), signaling that unique functional fingerprints of the experts have been blurred into a redundant subspace. In such a scenario, the router lacks the discriminative basis to perform meaningful, specialized dispatching.
In contrast, our method explicitly resolves this misalignment by permuting neurons to form a functionally coherent coordinate system prior to merging. Figure 3(c) demonstrates that this alignment dramatically restores inter-expert specialization, with CKA scores returning to near-optimal levels comparable to those of the original experts (Figure 3(a)). This provides empirical evidence that our alignment strategy effectively mitigates parameter space misalignment.
4.5 Analysis of Anchor Model Selection
Figure 4: Performance analysis of Symphony-MoE. (a) Average in-distribution (ID) and out-of-distribution (OOD) scores under different anchor model choices. (b) Impact of increasing the number of experts from 1 to 4 on ID and OOD performance.
[…]
As shown in the Figure 4(a), using the general-purpose expert model $M_{1}$ as the anchor yields the highest average scores on both ID and OOD tasks. This suggests that a balanced and non-specialized functional space plays a critical role in constructing an effective “omnipotent” MoE. It facilitates collaboration among experts without constraining them to overly narrow subspaces.
4.6 Analysis of the Number of Experts
[…]
4.7 Ablation Analysis
Model / Configuration ID Avg. OOD
(+) Full Method 44.07 35.26
(-) No Functional Alignment 33.94 26.92
(-) Naive Attention Merge 37.28 31.77
(-) Naive Embedding Merge 40.19 31.95
(-) Biased Calibration Data 39.70 29.44
Table 3: Ablation analysis of key components in Symphony-MoE (1.5B x 4).
We assessed the impact of key components on the performance of the Symphony-MoE model (1.5B × 4 configuration) through an ablation study, where each core module was individually removed or replaced. The results are presented in Table 3. Specifically, Naive Attention Merge replaces the SLERP fusion strategy with simple linear averaging when constructing the shared backbone for self-attention layers in Stage 1. This setting quantifies the contribution of SLERP to preserving the functional integrity of attention mechanisms. Naive Embedding Merge replaces MergeKit with a naive linear averaging of overlapping token embeddings during the construction of the shared embedding layer. Biased Calibration Data alters the sampling strategy of the calibration dataset $D_{\mathrm{cal}}$ in the alignment stage, using samples exclusively from the training set of the general expert $M_{1}$ rather than from all expert domains.
The results highlight the critical role of functional alignment: removing it causes catastrophic performance collapse, confirming that resolving parameter space inconsistencies is fundamental to the framework. Replacing SLERP with linear averaging leads to substantial degradation, supporting our hypothesis that the geometry of attention weights is essential for preserving functionality. While simplifying the embedding fusion yields only a moderate drop, it nonetheless indicates the value of layer-aware merging in the shared backbone. Notably, using biased calibration data results in the most severe performance decline after removing alignment, underscoring the importance of domain diversity in constructing $D_{\mathrm{cal}}$ for effective inter-expert coordination.
5 Conclusion
This paper presents Symphony-MoE, a framework for building powerful Mixture-of-Experts (MoE) models by upcycling experts from multiple, diverse pre-trained sources. We tackle the core challenge of parameter space misalignment through a two-stage process: training-free functional alignment and lightweight router training. Experiments show that Symphony-MoE consistently outperforms baselines on multi-domain tasks and generalizes well to out-of-distribution data. This approach offers a scalable path to leverage the collective knowledge embedded in existing specialized models—beyond the limitations of single-model upcycling.
Limitations and future work. Our current framework requires that all source models share an identical architecture. This constraint limits the pool of potential experts that can be integrated. Future work could focus on developing more advanced alignment techniques capable of harmonizing models with minor architectural dissimilarities, such as different activation functions or normalization strategies.
The quality of the functional alignment is contingent on the diversity and representativeness of the calibration dataset. While we proposed a simple and effective sampling strategy, optimizing the curation for a highly diverse set of experts is a non-trivial challenge.

### v2

@@
 Mixture-of-Experts (MoE) models enable scalable performance by activating large parameter sets sparsely, minimizing computational overhead.
-To circumvent the prohibitive cost of training MoEs from scratch, recent work employs upcycling, reusing a single pre-trained dense model by replicating its feed-forward network (FFN) layers into experts.
+To mitigate the prohibitive cost of training MoEs from scratch, recent work employs upcycling, reusing a single pre-trained dense model by replicating its feed-forward network (FFN) layers into experts.
 However, this limits expert diversity, as all experts originate from a single pre-trained dense model.
-This paper addresses this limitation by constructing powerful MoE models using experts sourced from multiple identically-architected but disparate pre-trained models (e.g., Llama2-Chat and Code Llama).
+This paper addresses this limitation by constructing powerful MoE models using experts sourced from multiple identically-architected but disparate pre-trained models (e.g., Qwen2.5-Coder and Qwen2).
 A key challenge lies in the fact that these source models occupy disparate, dissonant regions of the parameter space, making direct upcycling prone to severe performance degradation.
@@
 First, we establish this harmony in a training-free manner: we construct a shared backbone via a layer-aware fusion strategy and, crucially, alleviate parameter misalignment among experts using activation-based functional alignment.
-Subsequently, a single lightweight stage of router training coordinates the entire architecture.
+Subsequently, a stage of post-training coordinates the entire architecture.
 Experiments demonstrate that our method successfully integrates experts from heterogeneous sources, achieving an MoE model that significantly surpasses baselines in multi-domain tasks and out-of-distribution generalization.
@@
 2024) share a fundamental limitation: reliance on a single pre-trained checkpoint (see Figure 1).
-Consequently, all experts originate from identical parameter space, regardless of subsequent specialization.
+Consequently, all experts originate from an identical parameter space, regardless of subsequent specialization.
 This “single-origin” constraint intrinsically caps expert diversity.
@@
 In Stage 1, disparate source models are harmonized without training: non-FFN layers are merged into a shared backbone using techniques such as SLERP, while FFN layers are aligned through activation-based neuron permutation.
-In Stage 2, only the router is trained while keeping the backbone and expert parameters frozen, enabling coordination among the now-compatible experts.
+In Stage 2, the experts, router, and shared backbone undergo post-training to enable coordination among the now-compatible components.
 Thus, this paper addresses a core research question: Can we effectively assemble MoE models by upcycling experts from disparately initialized source models?
@@
 This training-free process yields a pool of diverse yet functionally compatible experts, setting the stage for efficient coordination.
-The second stage implements lightweight global coordination.
-We freeze the expert layers and shared backbone to train the router independently.
-To prevent expert collapse and ensure balanced utilization, this stage critically employs an auxiliary loss function promoting load balancing across experts.
+The second stage performs global coordination through post-training.
+To prevent expert collapse and encourage balanced expert utilization, this stage introduces a load-balancing loss that promotes uniform routing across experts.
 The core contributions of this work are threefold:
-(1) We investigate the problem of upcycling multiple identically architected yet independently pretrained models into a unified Mixture-of-Experts (MoE) system.
-We identify parameter space misalignment as a key barrier to this process..
-(2) We introduce a lightweight framework that addresses expert incompatibility without gradient-based training, combining layer-aware fusion and activation-driven functional alignment to preserve the specialized knowledge embedded in each source model.
-(3) Through comprehensive experiments on in-distribution and out-of-distribution datasets, we show that our method achieves strong performance and generalizes well to out-of-distribution settings, demonstrating the feasibility of upcycling disparate pre-trained models into unified expert systems.
+(1) We investigate the problem of upcycling multiple identically architected yet independently pretrained models into a unified MoE model.
+We identify parameter space misalignment as a key barrier to this process.
+(2) We introduce a harmonization framework that addresses expert incompatibility, combining layer-aware fusion and activation-driven functional alignment to preserve the specialized knowledge embedded in each source model.
+(3) Through comprehensive experiments on in-distribution and out-of-distribution datasets, we show that our method achieves strong performance and generalizes well to out-of-distribution settings, demonstrating the feasibility of harmonizing disparate pre-trained models into a coherent MoE.
 2 Related Works
@@
 2022; He et al.
+2024; Zhao et al.
 2024), involves converting a pre-trained dense model into an MoE by replicating its feed-forward network (FFN) layers to form multiple, initially identical experts.
@@
 The challenge of combining disparate pre-trained models is the central focus of the model merging field (Tao et al.
-2024).
+2024; Imfeld et al.
+2023).
 To merge truly “stranger” models, advanced techniques explicitly resolve misalignment.
@@
 Our work bridges this gap.
-We integrate permutation-based alignment directly into the MoE construction pipeline, repurposing these techniques for a new architectural goal.
+We integrate permutation-based alignment into the MoE construction pipeline, repurposing these techniques for a new goal.
 We use alignment to make disparate experts functionally compatible.
@@
 The first stage operates entirely without training, resolving the critical parameter space misalignment problem through training-free functional alignment.
-The second stage then trains a router that coordinates the now-compatible experts, unlocking their collective capabilities.
+The second stage then coordinates the now-compatible experts, unlocking their collective capabilities.
 3.1 Preliminaries and Problem Formulation
@@
 The goal is to construct a single, powerful MoE model, $M_{MoE}$, by leveraging the FFN layers of the source models as its experts.
-The fundamental barrier to this goal is the severe parameter space misalignment between the models, where the parameters $\theta_{i}$ and $\theta_{j}$ (for $i\neq j$) occupy incompatible numerical and semantic spaces.
+The fundamental barrier to this goal is the severe parameter space misalignment between the models, where the parameters $\theta_{i}$ and $\theta_{j}$ (for $i\neq j$) occupy incompatible semantic spaces.
 3.2 Stage 1: Training-Free Functional Alignment
-Our first stage assembles a cohesive MoE model from diverse parent models entirely without training.
-This process involves two key steps: fusing the shared architectural backbone and functionally aligning the FFN layers, which will serve as the experts.
+Our first stage assembles a MoE model from dense models entirely without training.
+This process involves two key steps: fusing the shared architectural backbone and functionally aligning the FFN expert layers.
 Shared Backbone Construction.
@@
 1.
-Activation Collection: We construct a small and diverse calibration set $D_{\mathrm{cal}}$ by sampling 5% from each instruction fine-tuning dataset $D_{1},D_{2},\ldots,D_{N}$} used in dense model training.
+Activation Collection: We construct a small and diverse calibration set $D_{\mathrm{cal}}$ by sampling equally from each continue pre-training dataset $D_{1},D_{2},\ldots,D_{N}$} used in dense model training.
 Each instruction in $D_{\mathrm{cal}}$ is passed through both the anchor model $M_{1}$ and a target model $M_{i}$ to extract their FFN activation matrices at layer $l$, denoted as $A_{1}^{(l)}$ and $A_{i}^{(l)}$, respectively.
+We sample activations using a tokenizer with a merged vocabulary, extract the post-nonlinearity activation values, and perform normalization.
 2. Permutation Matching: We find the optimal permutation matrix $P_{i}^{(l)}$ that aligns the neurons of $M_{i}$ to $M_{1}$ by solving the linear assignment problem:
@@
 2025).
-This involves permuting the output dimension of the first linear layer’s weights $W_{up}$ and the input dimension of the second’s $W_{down}$ to maintain functional equivalence:
-$W^{\prime}_{up,i}=W_{up,i}^{(l)}P_{i}^{(l)},\quad W^{\prime}_{down,i}=(P_{i}^{(l)})^{T}W_{down,i}^{(l)}$ (3)
+This involves permuting the output dimension of the first linear layer’s weights $W_{up}$, $W_{gate}$, and the input dimension of the second linear layer’s $W_{down}$ to maintain harmonization:
+$W^{\prime}_{up,i}=W_{up,i}^{(l)}P_{i}^{(l)},\quad W^{\prime}_{gate,i}=W_{gate,i}^{(l)}P_{i}^{(l)}$ (3)
+$W^{\prime}_{down,i}=(P_{i}^{(l)})^{T}W_{down,i}^{(l)}$ (4)
 This procedure results in $N$ functionally aligned experts for each layer, all operating in a consistent parameter space.
-We analyzed the computational cost to verify the scalability. The analysis can be found in the Appendix.
-3.3 Stage 2: Activating the MoE via Lightweight Router Training
+We analyzed the computational cost to verify the scalability. The time complexity of this solution is polynomial. The analysis can be found in the Appendix.
+3.3 Stage 2: Activating the MoE via Post-training
 In Stage 2, we integrate the constructed shared backbone with the expert models and introduce a randomly initialized router, implemented as a simple linear layer.
-The expert layers and shared backbone—which encode knowledge from the source models—are frozen, and only the router is trained.
-The calibration dataset $D_{\mathrm{cal}}$, introduced in Stage 1, is reused for this purpose, enabling the router to learn how to coordinate the experts.
-We adopt a top-2 routing mechanism and train the router using a composite loss function:
+The calibration dataset $D_{\mathrm{cal}}$, introduced in Stage 1, is extended, enabling the MoE to learn how to coordinate the experts.
+The expert layers, shared backbone, and the router are trained on the extended calibration dataset $D_{cal}$.
+We adopt a top-2 routing mechanism and train the MoE model using a composite loss function:
 $\mathcal{L}_{total}=\mathcal{L}_{lm}+\lambda\cdot\mathcal{L}_{bal}$ (5)
-where $\mathcal{L}_{lm}$ is the standard causal language modeling objective, and $\mathcal{L}_{bal}$ is a load-balancing regularization term.
+where $\mathcal{L}_{lm}$ is the standard causal language modeling objective, and $\mathcal{L}_{bal}$ is a load-balancing regularization term introduced in Switch Transformer (Fedus, Zoph, and Shazeer 2022).
 The coefficient $\lambda$ is fixed at 0.01 in all experiments.
-$\mathcal{L}_{bal}$ encourages uniform expert utilization, which mitigates expert collapse and allows the model to fully leverage the diversity of its specialists.
-Notably, the router is not provided with explicit supervision indicating which expert to select for each token, as there is no ground-truth mapping between tokens and experts.
-Instead, it learns expert assignment implicitly through training on instruction–response pairs in $D_{\mathrm{cal}}$.
+Notably, the router is not provided with explicit supervision indicating which expert to select for each token.
+It learns expert assignment implicitly through continual pre-training on $D_{\mathrm{cal}}$.
 4 Experiments
@@
 ID OOD
-MMLU GSM8K BBH HumanEval TruthfulQA Avg. MedCQA
+MMLU GSM8K BBH HumanEval TruthfulQA Avg.* MedCQA
 Dense (0.5B)
@@
 “ANC” stands for anchor model.
+“Avg*” stands for the average scores of ID datasets.
 ID OOD
@@
 “ANC” stands for anchor model.
+“Avg*” stands for the average scores of ID datasets.
 To ensure a rigorous evaluation of our proposed merging method, we first constructed a set of four disparately initialized dense models at scale of 0.5/1.5B.
@@
 2021), BBH (General, Reasoning)  (Suzgun et al.
-2022), HumanEval (Code) (Chen et al.
-2021), and TruthfulQA (Science) (Lin, Hilton, and Evans 2022).
+2022), HumanEval (Code) (Chen 2021), and TruthfulQA (Science) (Lin, Hilton, and Evans 2022).
 We also include an out-of-distribution benchmark, MedCQA (Medicine) (Shoham and Rappoport 2024).
@@
 This misalignment leads to representational collapse, erasing the distinct capabilities of individual experts.
-Our experiments were conducted on 8 NVIDIA V100 GPUs.
-The dense model was implemented using LLaMA-Factory (Zheng et al.
-2024), and trained for for 2 epoches.
-Detailed training configurations are provided in the Appendix.
-In Stage 1: Alignment, we randomly sampled from each training dataset to construct the calibration set $D_{\mathrm{cal}}$, resulting in a total of 10.404 million tokens.
+Our experiments were conducted on 24 NVIDIA V100 GPUs. The dense model was implemented using LLaMA-Factory (Zheng et al. 2024), and trained for 2 epochs.
+In Stage 1: Alignment, we randomly sampled from each training dataset to construct the calibration set $D_{\mathrm{cal}}$, resulting in a total of 10.4M tokens.
+The training data is sampled from SlimPajama [^1] https://huggingface.co/datasets/cerebras/SlimPajama-627B (General, Code), Finemath [^2] https://huggingface.co/datasets/HuggingFaceTB/finemath, and Scientific Papers [^3] https://huggingface.co/datasets/scientific˙papers.
 To ensure consistent alignment and integration across expert models, we fixed a general-purpose expert model $M_{1}$ as the anchor model.
 To ensure that the knowledge of each expert is equally integrated and to avoid introducing any bias, the weights for merging the 4 dense models are set to 0.25.
-In Stage 2: Router Training, we set the cutoff length to 2048 and train the model for 6 epochs with a batch size of 2.
+In Stage 2: Post-training, we increased the sampled data size, expanding $D_{\mathrm{cal}}$ to 5B tokens.
+We set the cutoff length to 2048 and train the model for 6 epochs with a batch size of 2.
 The learning rate is fixed at 5e-5, and optimization is performed using AdamW (Zhuang et al.
@@
 A lower CKA score indicates greater functional specialization.
-We evaluate three scenarios: (a) the original, unmerged expert models as a reference baseline; (b) a naive merging approach without alignment; and (c) our proposed method employing activation-based functional alignment.
+We evaluate three scenarios: (a) the original, unmerged expert models as a reference baseline; (b) a naive merging approach without alignment; (c) our method employing activation-based functional alignment.
 As shown in Figure 3, the naive merging strategy leads to a severe representational collapse.
@@
 This results in inflated inter-expert CKA scores (e.g., 0.65–0.75), signaling that unique functional fingerprints of the experts have been blurred into a redundant subspace.
-In such a scenario, the router lacks the discriminative basis to perform meaningful, specialized dispatching.
+In such a scenario, the router lacks the discriminative basis to perform specialized dispatching.
 In contrast, our method explicitly resolves this misalignment by permuting neurons to form a functionally coherent coordinate system prior to merging.
 Figure 3(c) demonstrates that this alignment dramatically restores inter-expert specialization, with CKA scores returning to near-optimal levels comparable to those of the original experts (Figure 3(a)).
-This provides empirical evidence that our alignment strategy effectively mitigates parameter space misalignment.
+This provides empirical evidence that our harmonization framework effectively mitigates parameter space misalignment.
 4.5 Analysis of Anchor Model Selection
@@
 (-) No Functional Alignment 33.94 26.92
+   (-) Align 80% of the Neurons 43.57 34.61
 (-) Naive Attention Merge 37.28 31.77
@@
 The results highlight the critical role of functional alignment: removing it causes catastrophic performance collapse, confirming that resolving parameter space inconsistencies is fundamental to the framework.
+Aligning only 80% of the neurons did not result in a significant decrease in the model’s performance, demonstrating the robustness of the alignment method.
 Replacing SLERP with linear averaging leads to substantial degradation, supporting our hypothesis that the geometry of attention weights is essential for preserving functionality.
@@
 5 Conclusion
-This paper presents Symphony-MoE, a framework for building powerful Mixture-of-Experts (MoE) models by upcycling experts from multiple, diverse pre-trained sources.
-We tackle the core challenge of parameter space misalignment through a two-stage process: training-free functional alignment and lightweight router training.
-Experiments show that Symphony-MoE consistently outperforms baselines on multi-domain tasks and generalizes well to out-of-distribution data.
-This approach offers a scalable path to leverage the collective knowledge embedded in existing specialized models—beyond the limitations of single-model upcycling.
-Limitations and future work.
-Our current framework requires that all source models share an identical architecture.
+This paper presents Symphony-MoE, a framework for building powerful MoE models by upcycling experts from multiple, diverse pre-trained sources.
+We tackle the core challenge of parameter space misalignment through a two-stage process.
+Experiments show that Symphony-MoE outperforms baselines on multi-domain tasks and generalizes well to out-of-distribution data.
+This approach offers a scalable path to leverage the collective knowledge embedded in existing specialized models.
+Limitations.
+Our current framework requires that all source models share an identical architecture and are all language models.
 This constraint limits the pool of potential experts that can be integrated.
-Future work could focus on developing more advanced alignment techniques capable of harmonizing models with minor architectural dissimilarities, such as different activation functions or normalization strategies.
-The quality of the functional alignment is contingent on the diversity and representativeness of the calibration dataset.
-While we proposed a simple and effective sampling strategy, optimizing the curation for a highly diverse set of experts is a non-trivial challenge.
+Future work could focus on developing more advanced alignment techniques capable of harmonizing models with minor architectural dissimilarities or modality differences.
+The effectiveness of functional alignment depends on the diversity of the calibration dataset.
+Although we propose a straightforward sampling strategy, optimizing the process to ensure a highly diverse set of experts presents a significant challenge.
+Appendix A Computational Cost Analysis of Pairwise Inter-Expert Alignment
+The scalability of the Symphony-MoE framework is critically dependent on the computational cost of its alignment stage.
+This section provides an analysis of the computational complexity of the Pairwise Inter-Expert Alignment process, demonstrating its efficiency and viability for large-scale models.
+The alignment process consists of three main steps: 1) Activation Collection, 2) Permutation Matching, and 3) Weight Remapping.
+We analyze the cost of each step for a single layer, which can then be multiplied by the number of layers to determine the total cost.
+1.
+Activation Collection: This step involves performing a forward pass through the anchor model ($M_{1}$) and each non-anchor model ($M_{i}$) to extract FFN activation matrices.
+The cost is proportional to the size of the calibration dataset ($D_{cal}$) and the number of parameters in the models.
+Let $C_{forward}$ be the cost of a single forward pass.
+For $N$ models, this step is performed for all N1N-1 non-anchor models.
+Thus, the total cost for this step is $(N-1)\times C_{forward}$.
+Since the forward pass is a standard operation in neural networks, this cost is manageable.
+2.
+Permutation Matching: For each layer, we solve a linear assignment problem to find the optimal permutation matrix ($P_{i}^{(l)}$) that aligns the neurons of a non-anchor model to the anchor model.
+This is formulated as:
+$\min_{P\in\mathcal{P}}\|A_{1}^{(l)}-A_{i}^{(l)}P\|_{F}^{2}$
+where $\mathcal{P}$ is the set of permutation matrices.
+This problem is efficiently solved using the Hungarian algorithm.
+The computational complexity of the Hungarian algorithm for a $d\times d$ matrix, where $d$ is the number of neurons in the FFN layer, is $O(d^{3})$.
+3.
+Weight Remapping: The computed permutation matrix is then applied to the weights of the FFN layers.
+This involves permuting the columns of the first linear layer’s weights ($W_{up,i}^{(l)}$) and the rows of the second linear layer’s weights ($W_{down,i}^{(l)}$).
+$W_{up,i}^{\prime}=W_{up,i}^{(l)}P_{i}^{(l)}$
+$W_{down,i}^{\prime}=(P_{i}^{(l)})^{T}W_{down,i}^{(l)}$
+The cost of these matrix multiplications for an FFN layer with input dimension $d_{in}$, hidden dimension $d$, and output dimension $d_{out}$ is $O(d_{in}\times d\times d)$ for the first layer and $O(d\times d\times d_{out})$ for the second.
+In typical transformer architectures, $d_{in}=d_{out}$, so the total cost for remapping is approximately $O(d^{3})$.
+Overall Scalability: The dominant computational cost in the alignment of a single layer is the permutation matching and weight remapping, both of which are polynomial in the number of neurons ($d$).
+The total cost for aligning all FFN layers in N1N-1 non-anchor models is proportional to $(N-1)\times L\times O(d^{3})$, where $L$ is the number of layers.
+While the cubic complexity with respect to the number of neurons may seem significant, in practice, this alignment is a one-time, offline process performed before post-training.
+The cost is independent of the size of the training dataset for the final MoE model and is parallelizable across layers and models, making it a scalable solution for constructing large-scale MoE models from disparate pre-trained models.
+Appendix B Experimental Setup Details
+To create a set of specialized experts, we fine-tuned four distinct dense models.
+The foundation models include Qwen2.5-Base, its predecessor Qwen2-Base, and the code-specialized Qwen2.5-Coder.
+This selection was made to introduce disparity across model versions, pre-training data, and downstream tasks.
+The instruction fine-tuning for all dense models was performed for 2 epochs with a consistent set of hyperparameters to ensure comparability.
+For the model’s training configuration, we selected the AdamW optimizer with a learning rate set to 5e-5.
+During training, a batch size of 2 was used, and to prevent exploding gradients, we employed gradient clipping, limiting the max gradient norm to 1.0.
+The sequence cutoff length was set to 2048.
+Additionally, for regularization, a weight decay of 0.01 was applied, and a Cosine Decay scheduler was used to adjust the learning rate dynamically.
+The calibration dataset $D_{cal}$, used in Stage 1, was constructed by sampling equally from each of the instruction fine-tuning datasets (See Table 5), resulting in a total of 10.404M tokens.
+This ensures a balanced representation of each expert’s domain during alignment and router training.
+For the post-training in Stage 2, we trained for 6 epochs using the AdamW optimizer with a learning rate of 5e-5.
+Domain Datasets for Source Model Construction
+General Alpaca
+Math MetaMathQA
+Code CodeAlpaca
+Science SciQAG
+Table 4: Explanation of instruction-finetuning datasets for source models. This batch of data was entirely used for training and was not sampled.
+Domain Source of the Calibration Dataset
+General SlimPajama-Wikipedia
+Math Finemath
+Code SlimPajama-Github
+Science Scientific-Papers
+Table 5: Explanation of the source of the Calibration Dataset $D_{cal}$.
+The number of tokens sampled from each data source is the same to ensure a balanced composition of $D_{cal}$.
+The random seed for sampling is set to 114514.
+Appendix C Description and Reproduction Details of Baselines
+MMLU GSM8K BBH HumanEval TruthfulQA Avg. MedCQA
+Dense (1B)
+General ($M_{1}$, ANC) 61.35±0.39 35.14±1.09 45.05±0.55 40.15±3.80 30.11±1.60 42.36 31.55±0.20
+Math ($M_{2}$) 58.10±0.39 38.95±1.33 47.05±0.54 39.50±3.79 32.05±1.62 43.13 31.89±0.20
+Code ($M_{3}$) 45.20±0.41 37.10±1.33 38.22±0.55 46.88±3.81 27.85±1.51 39.05 29.81±0.19
+Science ($M_{4}$) 57.90±0.40 34.20±1.30 37.95±0.53 38.14±3.74 33.05±1.57 40.25 30.98±0.19
+MoE (1B $\times$ 4)
+BTX 47.33±0.88 32.15±0.45 41.80±0.58 31.77±3.15 26.95±1.15 36.00 28.15±0.18
+BAM 52.95±0.75 38.10±0.61 46.85±0.59 40.10±2.89 28.88±1.89 41.38 30.85±0.18
+Drop 59.20±0.81 37.85±1.12 46.90±0.43 42.15±1.89 32.95±1.49 43.81 34.25±0.19
+Symphony (Ours) 60.88±0.31 38.80±1.09 47.11±0.36 45.95±3.69 33.15±2.31 47.18 37.33±0.22
+Table 6: Performance comparison of dense (Llama 3.2 1B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data.
+All baselines in our comparison follow a common two-stage pipeline: upcycling → post-training.
+In the post-training stage, to ensure a fair comparison, all models are trained solely on the extended calibration dataset $D_{\mathrm{cal}}$ (as constructed in our framework), using the AdamW optimizer for 6 epochs with a learning rate of 5e-5.
+In the post-training phase, all modules are trainable, and the learning rate remains consistent.
+The upcycling strategies adopted by each baseline are as follows:
+(a) BTX constructs experts by directly reusing the FFN weights from each dense model. The shared backbone is formed by linearly averaging all model weights.
+(b) BAM reuses FFN weights and partial attention weights ($W^{q}$, $W^{o}$) to construct experts, while the remaining weights are linearly averaged to form the shared backbone.
+(c) Drop-Upcycling reuses FFN weights and applies Gaussian perturbation to randomly selected parameters to prevent expert homogenization.
+The remaining weights are reused and averaged to build the shared backbone, and the expert modules are also updated during training.
+Appendix D Additional Experimental Results
+For small-scale models, the Llama series uses Llama3.2 1B for all versions ($M_{1},M_{2},M_{3},M_{4}$) to isolate the impact of varying training tasks, data, and parameters on model performance.